OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

y = 4*x+11

OpenStudy (john_es):

The above solution is correct. Do you find it in your answers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (john_es):

Please post them, and we can compare. May be it is given in point slope form. It is the second option.

OpenStudy (anonymous):

i posted what do you mean

OpenStudy (john_es):

The second option is the correct. You can check it, \[y-y_0=m(x-x_0)\] m is the slope. x0 and y0 are the coordinates of the point given ;).

OpenStudy (anonymous):

thanks! can you help me on more?

OpenStudy (john_es):

Yes.

OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (john_es):

If it is parallel, it must have the same slope -3. And if pass through (2,6) then \[y=-3x+n\] \[6=-3\cdot2+n\Rightarrow n=12\]So, \[y=-3x+12\]

OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (john_es):

Sorry, it is the first one.

OpenStudy (anonymous):

one more!! Which of the following is the slope between the points (0, -7) and (4, 12)? 19 fourths -19 fourths 4 nineteenths -4 nineteenths

OpenStudy (john_es):

The first one.

OpenStudy (anonymous):

thank you john you been a big help

OpenStudy (john_es):

;)